OpenStudy (anonymous):

can someone please help me with this problem, I am really lost. x^4-32x^2+31.

OpenStudy (ribhu):

x^4-31x^2-x^2+31=0

OpenStudy (ribhu):

x^2(x^2-31)-1(x^2-31)=0

OpenStudy (anonymous):

It said I had to factor out the polynomial.

OpenStudy (ribhu):

(x^2-1)(x^2-31)=0

OpenStudy (ribhu):

this is what i did for u

OpenStudy (anonymous):

Ok thank you.